# Better Call SAL: Towards Learning to Segment Anything in Lidar

### v1

@@ Abstract
We propose the SAL (Segment Anything in Lidar) method consisting of a text-promptable zero-shot model for segmenting and classifying any object in Lidar, and a pseudo-labeling engine that facilitates model training without manual supervision. While the established paradigm for Lidar Panoptic Segmentation (LPS) relies on manual supervision for a handful of object classes defined a priori, we utilize 2D vision foundation models to generate 3D supervision “for free”. Our pseudo-labels consist of instance masks and corresponding CLIP tokens, which we lift to Lidar using calibrated multi-modal data. By training our model on these labels, we distill the 2D foundation models into our Lidar SAL model. Even without manual labels, our model achieves $91\%$ in terms of class-agnostic segmentation and $44\%$ in terms of zero-shot LPS of the fully supervised state-of-the-art. Furthermore, we outperform several baselines that do not distill but only lift image features to 3D. More importantly, we demonstrate that SAL supports arbitrary class prompts, can be easily extended to new datasets, and shows significant potential to improve with increasing amounts of self-labeled data.
@@
The appendix provides additional implementation and design details of the SAL pseudo-label engine (A.1) and zero-shot model (A.2), along with a discussion on different training strategies with partial pseudo-labels (A.4). In A.3, we discuss the zero-shot text prompt engineering. Further ablations on our pseudo-labels and model are provided in B and C, respectively, including fine-grained, per-class results (D) and detailed statistics on the generated pseudo-labels (B.1) and time and compute analysis for pseudo-labeling (B.2). We conclude with a discussion on detailed qualitative results in E.
@@
Figure 1: The SAL (Segment Anything in Lidar) model performs class-agnostic instance segmentation (i) and zero-shot classification via text prompting. This allows us to not only predict panoptic segmentation (ii) for fixed class vocabularies but segment any object (iii and iv) in a given Lidar scan.
@@ 1 Introduction
Towards segmenting anything in Lidar.  To address these challenges, we propose SAL, which consists of a text-promptable zero-shot model (Fig.˜2, left) for panoptic segmentation of arbitrary objects (Fig.˜1), and pseudo-labeling engine (Fig.˜2, right) that facilitates model training directly from raw sensory data. The SAL pseudo-label engine automatically labels Lidar sequences using image segmentation [31] and vision-language models [67]. We utilize SAM [31] to generate class-agnostic masks in images, and CLIP [67] to generate per-mask tokens that connect visual features to language, and finally, transfer both to Lidar using a calibrated sensory setup. Even though generated pseudo-labels only partially cover Lidar scans and are inherently noisy due to errors in the image-level generation process and imperfect sensory calibration, we demonstrate their effectiveness as self-supervised training data for the SAL zero-shot model. In contrast to prior work in LPS, SAL is the first model that can be prompted with any new vocabulary without re-training. Different from recent advances [58] that lift image features to 3D without distillation into a model, SAL does not require image features during inference and segments full $360^{\circ}$ point clouds.
Talk to your SAL Our model predicts Lidar segmentation masks and their corresponding CLIP tokens, which allow us to perform zero-shot classification of segmented objects using arbitrary text prompts (Fig.˜1). We evaluate SAL by prompting our zero-shot model on standard benchmarks for LPS. For class-agnostic segmentation, we reach $91\%$ of the performance of manually supervised baselines. By classifying segmented objects in a zero-shot manner, we report the first (very encouraging) results for zero-shot LPS and reach 4040% and $44\%$ of the baselines on SemanticKITTI and nuScenes, respectively. Beyond that, as shown in Fig.˜1, we can prompt SAL model to segment objects outside of existing Lidar dataset class vocabularies.
@@ 3.4 SAL Zero-Shot Model
Universal architecture.  Instead of relying on highly specialized and engineered LPS models, we base the SAL model on a universal Transformer decoder architecture (Fig.˜3(b)), similar to [45]. Its 2D counterparts [9, 11] provide top-performing results across semantic, instance, and panoptic segmentation for images. With increasing amounts of (pseudo-)labeled data, Transformer decoders have the potential to achieve a similar impact in the Lidar domain. Following [45], we deploy a Minkowski U-Net [12] backbone for feature extraction followed by a Transformer decoder architecture with object query to point/voxel feature cross-attention. Training our model to segment and classify any object and perform in a zero-shot evaluation setting requires a unique design of the final task heads.
@@
How to train with partial labels?  Partially (pseudo-)labeled point clouds present a challenge for SAL model training. For example, only $14\%$ of points are pseudo-labeled due to low camera coverage in SemanticKITTI (see Tab.˜B.1). If we naively train SAL on partially (pseudo-)labeled point clouds, the objectness loss $\mathcal{L}_{obj}$ would penalize any segmentation in these regions as a false positive, thereby teaching the model to ignore these points entirely. Empirically, we determined that the most effective training strategy is to (i) remove unlabeled points from point clouds, (ii) utilize standard data augmentations (rotations, flipping, scaling, and translations), in conjunction with proposed FrankenFrustum augmentation (detailed in Sec.˜A.4) that mimics fully-labeled point clouds during training by randomly removing unlabeled points and replicating labeled frustum regions around the vertical axis. This augmentation does not increase the overall label coverage, but our ablations in Sec.˜4.2 show its effectiveness in reducing the domain gap between training and inference input clouds.
@@ 4.1 Experimental Setup
We evaluate SAL on two public Lidar Panoptic Segmentation (LPS) datasets, SemanticKITTI [6, 7] and nuScenes [18]. We utilize provided, human-annotated ground-truth (GT) panoptic segmentation labels solely for evaluation purposes, linear probing ablations, and baselines. We train SAL model only using our self-supervised pseudo-labels (Sec.˜3.3). To demonstrate the general applicability of SAL to other domains and datasets, we show qualitative results of SAL trained on Waymo Open [75] dataset, which contains no GT LPS labels.
@@ 4.2 SAL Ablations
Learning with partial labels.  For the single-camera setup [6, 7], our pseudo-labels only cover $14\%$ of the input point clouds (see Tab.˜B.1). Training on partial label coverage presents a challenge for our model. In Tab.˜2, we only evaluate class-agnostic segmentation by applying SO and SM (see Sec.˜4.1). We compare SAL models trained using pseudo and GT labels. The Frustum Filter removes all unlabeled points not visible in the camera frustum during training and/or evaluation. As shown in Tab.˜2, removing unlabeled points from pseudo-labeled point clouds during training has a significant benefit on the performance ($3^{rd}$ row, $59.3$ PQ) as compared to the variant, where we merely ignore unlabeled region during the training ($2^{nd}$ row, $22.2$ PQ). This is likely due to data imbalance in a single-camera setup, leading to a class imbalance between labeled/unlabeled points, incentivizing SAL to suppress predictions.
While utilizing standard augmentations (rotations, flipping, scaling, and translations) is crucial, we observe that performing our FrankenFrustum augmentation, which concatenates the visible portion of the point cloud around the z-axis (see Sec.˜3.4), significantly improves results ($4^{th}$ row, $62.5$ PQ). Mixing point clouds from different scans ($5^{th}$ row, $62.8$ PQ) further boosts results. Overall, we obtain $69$ PQ from GT supervision ($1st$ row) and $62.8$ PQ when using our pseudo-labels ($5th$ row). Remarkably, our labels yield $91\%$ of the GT performance while covering only $14\%$ of the input point cloud (see Tab.˜B.1). To further contextualize these results, we note that this evaluation is performed only on classes for which GT instance labels are available. As shown in Fig.˜1, our model learns to segment a much larger variety of classes. Evaluating only the subset of points visible in the camera ($14\%$ of all points, rows 6&76\&7), the gap between the GT-supervised model ($71.8$ PQ) and pseudo-supervised model ($70.7$ PQ) shrinks further. We conclude that by training SAL model using pseudo-labels, we distill the notion of objectness from SAM [31] into our model and obtain segmentation capabilities, similar to the model trained on GT data.
@@
Table 2: Scaling. In contrast to pre-defined GT data, increasing the number of queries SAL improves performance on stuff classes. By increasing the amount of labeled data, we further improve segmentation performance.
@@
$\times$ $23.6$ $31.2$ $67.0$ $29.0$
Semantic distillation.  At inference, SAL predicts a binary segmentation mask and objectness score for each query. To perform zero-shot classification, our model additionally predicts CLIP image tokens. This allows us to not rely on image features during inference but prompt our predicted tokens with arbitrary texts, i.e., object classes. Does the token head loss $\mathcal{L}_{token}$ successfully distill a notion of object semantics from CLIP into our model? To verify this, we perform linear probing experiments: we replace the token distillation head of a trained model with a linear classifier (backbone and instance decoder remain frozen) and train it using GT labels. As can be seen in Tab.˜3, distilling CLIP features into our model significantly improves linear probing performance ($33.1$ PQ, $40.0$ mIoU) compared to the baseline, not trained with the semantic distillation loss $\mathcal{L}_{token}$ ($20.0$ PQ, $23.4$ mIoU). This confirms that by distilling CLIP features, we inject a notion of semantics into our model, even though our model is not explicitly supervised with any labels containing semantic information – the notion of semantics is learned implicitly via CLIP feature distillation. Without this loss, our zero-shot model ($23.6$ PQ and $29.0$ mIoU) outperforms the linearly probed model tuned using GT semantic labels.
@@ 4.3 Class-Agnostic and Zero-Shot Lidar Panoptic Segmentation
Do we need a vision segmentation foundation model?  To generate a set with a sufficiently high recall, we need to cluster points with varying DBSCAN radius parameters (details in Sec.˜A.1). This leads to a large set of predicted masks per scan ($5,413$ avg.). While most of the correct segments are in this set, separating the signal from the noise is difficult. Prior works, such as [52], are limited to moving thing classes as they rely on motion cues to filter static point cloud regions and DBSCAN to group remaining points. Our core insight is that we can utilize vision foundation models that already learned a notion of objectness: SAM generates only $171$ masks per image, reduced to $45$ via flattening, and to $39$ masks after the unprojection, achieving a high recall with only a few instances, or in other words, a high signal-to-noise ratio.
@@
Zero-shot segmentation.  So far, we discussed class-agnostic segmentation. To obtain semantic class predictions per mask in a zero-shot fashion, we devise a simple baseline inspired by [58]. Once masks are unprojected from the image to Lidar space, we utilize associated CLIP features (directly extracted from images using [16]) to perform zero-shot classification via dot product between encoded class prompts and (lifted) CLIP features [67]. This baseline (SAM+DBS+CLIP) can be understood as an approach that directly lifts image features from state-of-the-art image-based models [16] to Lidar. We detail prompts in Sec.˜A.3.
@@
There is a significant performance drop between SAM+DBS (replace) baseline, evaluated with Semantic Oracle ($5^{th}$ row, $48.7$ PQ), and zero-shot baseline (SAM+DBS (replace)+CLIP), $7^{th}$ row, $17.9$ PQ. This is not surprising, as zero-shot Segmentation is a challenging problem in both image [16] and Lidar domains. To this end, even when transferring features of state-of-the-art image-based models to Lidar we observe a significant performance drop. Remarkably, SAL distills such noisy image-based features to a stronger model ($8^{th}$ row, $27.6$ PQ, +9.7+9.7 PQ). Finally, to emphasize the inherent limitation of baselines, relying on image features, we evaluate on full $360^{\circ}$ point clouds. SAL performs similarly in both settings. The baseline, on the other hand, relies on image features and is, therefore, unable to predict labels out of the camera frustum. This problem can be mitigated by utilizing a setup with denser camera coverage.
@@
Evaluation on super-classes.  We observe that the CLIP classifier often confuses related classes, e.g., car and other vehicle. We additionally evaluate our semantics by prompting with super-classes (as defined in [6]). This alone increases PQ to $57.5$ ($46.3$ when evaluating on full point clouds) and suggests significant potential for improving prompting and text-to-image alignment.
@@
Table 5: Lidar Panoptic Segmentation (LPS) on SemanticKITTI and nuScenes validation sets. We prompt our zero-shot SAL model with the respective class vocabularies and compare its performance to fully-supervised baselines. On SemanticKITTI and nuScenes, we reach $40\%$ and $46\%$ of the fully-supervised model, respectively. This gap reduces significantly when we evaluate super classes.
@@
SAL Zero-shot $23.6$ $31.2$ $67.0$ $21.7$ $25.1$ $29.0$ $46.3$ $60.1$ $76.1$ $52.3$
SAL Linear prob. $33.1$ $41.9$ $68.3$ $31.8$ $34.1$ $40.0$ $55.3$ $69.3$ $77.9$ $53.0$
nuScenes [18] PHNet [36] Full $74.7$ $84.2$ $88.2$ $74.0$ $75.9$ $79.7$
@@
SAL Zero-shot $31.3$ $39.8$ $74.3$ $38.8$ $23.8$ $28.7$ $42.1$ $53.2$ $74.7$ $41.0$
SAL Linear prob. $40.1$ $50.5$ $76.4$ $45.5$ $34.6$ $39.0$ $59.2$ $73.0$ $79.3$ $59.3$

### v2

@@ Abstract
We propose the SAL (Segment Anything in Lidar) method consisting of a text-promptable zero-shot model for segmenting and classifying any object in Lidar, and a pseudo-labeling engine that facilitates model training without manual supervision. While the established paradigm for Lidar Panoptic Segmentation (LPS) relies on manual supervision for a handful of object classes defined a priori, we utilize 2D vision foundation models to generate 3D supervision “for free”. Our pseudo-labels consist of instance masks and corresponding CLIP tokens, which we lift to Lidar using calibrated multi-modal data. By training our model on these labels, we distill the 2D foundation models into our Lidar SAL model. Even without manual labels, our model achieves $91\%$ in terms of class-agnostic segmentation and $54\%$ in terms of zero-shot LPS of the fully supervised state-of-the-art. Furthermore, we outperform several baselines that do not distill but only lift image features to 3D. More importantly, we demonstrate that SAL supports arbitrary class prompts, can be easily extended to new datasets, and shows significant potential to improve with increasing amounts of self-labeled data. We release all models and the code.
@@
The appendix provides additional implementation and design details of the SAL pseudo-label engine (0.A.1) and zero-shot model (0.A.2), along with a discussion on different training strategies with partial pseudo-labels (0.A.4). In 0.A.3, we discuss the zero-shot text prompt engineering. Further ablations on our pseudo-labels and model are provided in 0.B and 0.C, respectively, including fine-grained, per-class results (0.D) and detailed statistics on the generated pseudo-labels (0.B.1) and time and compute analysis for pseudo-labeling (0.B.2). We conclude with a discussion on detailed qualitative results in 0.E.
@@
Figure 1: The SAL (Segment Anything in Lidar) model performs class-agnostic instance segmentation (i) and zero-shot classification via text prompting. This allows us to not only predict panoptic segmentation (ii) for fixed class vocabularies but also segment any object (iii and iv) in a given Lidar scan.
@@ 1 Introduction
Towards segmenting anything in Lidar.  To address these challenges, we propose SAL, which consists of a text-promptable zero-shot model (Fig.˜2, left) for panoptic segmentation of arbitrary objects (Fig.˜1), and pseudo-labeling engine (Fig.˜2, right) that facilitates model training directly from raw sensory data. The SAL pseudo-label engine automatically labels Lidar sequences using image segmentation [31] and vision-language models [67]. We utilize SAM [31] to generate class-agnostic masks in images, and CLIP [67] to generate per-mask tokens that connect visual features to language, and finally, transfer both to Lidar using a calibrated sensory setup. Even though generated pseudo-labels only partially cover Lidar scans and are inherently noisy due to errors in the image-level generation process and imperfect sensory calibration, we demonstrate their effectiveness as self-supervised training data for the SAL zero-shot model. In contrast to prior work in LPS, SAL can be prompted with any semantic vocabulary during test-time without model re-training or tuning. Different from recent advances in zero-shot semantic segmentation [58] SAL does not require image features during inference and can segment full $360^{\circ}$ point clouds.
Talk to your SAL Our model predicts Lidar segmentation masks and their corresponding CLIP tokens, which allow us to perform zero-shot classification of segmented objects using arbitrary text prompts (Fig.˜1). We evaluate SAL by prompting our zero-shot model on standard benchmarks for LPS. For class-agnostic segmentation, we reach $91\%$ of the performance of manually supervised baselines. By classifying segmented objects in a zero-shot manner, we report the first (and very encouraging) results for zero-shot LPS and reach 4242% and $54\%$ of the supervised baselines trained on SemanticKITTI and nuScenes, respectively. Beyond that, as shown in Fig.˜1, we can prompt SAL model to segment objects outside of existing Lidar dataset class vocabularies.
@@ 3.4 SAL Zero-Shot Model
Universal architecture.  Instead of relying on highly specialized and engineered LPS models, we base the SAL model on a universal Transformer decoder architecture (Fig.˜3(b)), similar to [45]. Its 2D counterparts [9, 11] provide top-performing results across semantic, instance, and panoptic segmentation for images. With increasing amounts of (pseudo-)labeled data, Transformer decoders have the potential to achieve a similar impact in the Lidar domain. Following [45], we deploy a Minkowski U-Net [12] backbone for feature extraction followed by a Transformer decoder architecture with object query to point/voxel feature cross-attention. Training our model to segment and classify any object and perform the evaluation in a zero-shot setting requires a unique design of the final task heads.
@@
How to train with partial labels?  Partially (pseudo-)labeled point clouds present a challenge for SAL model training. For example, in SemanticKITTI, only $14\%$ of points are pseudo-labeled due to low camera coverage (see Tab.˜0.B.1). If we naively train SAL on partially (pseudo-)labeled point clouds, the objectness loss $\mathcal{L}_{obj}$ would penalize any segmentation in these regions as a false positive, thereby teaching the model to ignore these points entirely. Empirically, we determined that the most effective training strategy is to (i) remove unlabeled points from point clouds, (ii) utilize standard data augmentations (rotations, flipping, scaling, and translations), in conjunction with proposed FrankenFrustum augmentation (detailed in Sec.˜0.A.4) that mimics fully-labeled point clouds during training by randomly removing unlabeled points and replicating labeled frustum regions around the vertical axis. This augmentation does not increase the overall label coverage, but our ablations in Sec.˜4.2 show its effectiveness in reducing the domain gap between training and inference input clouds.
@@ 4.1 Experimental Setup
We evaluate SAL on two public Lidar Panoptic Segmentation (LPS) datasets, SemanticKITTI [6, 7] and nuScenes [18]. We utilize provided, human-annotated ground-truth (GT) panoptic segmentation labels solely for evaluation purposes, linear probing ablations, and baselines. We train SAL model only our pseudo-labels (Sec.˜3.3). To demonstrate the general applicability of SAL to other domains and datasets, we further show qualitative results of SAL trained on Waymo Open [75] dataset, which contains no GT LPS labels.
@@ 4.2 SAL Ablations
Learning with partial labels.  For the single-camera setup [6, 7], our pseudo-labels only cover $14\%$ of the input point clouds (see Tab.˜0.B.1). Training on partial label coverage presents a challenge for our model.
In Tab.˜2, we only evaluate class-agnostic segmentation by applying SO and SM (see Sec.˜4.1). We compare SAL models trained using pseudo and GT labels. The Frustum Filter removes all unlabeled points not visible in the camera frustum during training and/or evaluation. As shown in Tab.˜2, removing unlabeled points from pseudo-labeled point clouds during training shows a significant benefit on the performance ($3^{rd}$ row, $59.3$ PQ) as compared to the variant, where we merely ignore unlabeled region during the training ($2^{nd}$ row, $22.2$ PQ). This is likely due to data imbalance in a single-camera setup, leading to a class imbalance between labeled/unlabeled points, incentivizing SAL to suppress predictions.
While utilizing standard augmentations (rotations, flipping, scaling, and translations) is crucial, we observe that performing our FrankenFrustum augmentation, which concatenates the visible portion of the point cloud around the z-axis (see Sec.˜3.4), significantly improves results ($4^{th}$ row, $62.5$ PQ). Mixing point clouds from different scans ($5^{th}$ row, $62.8$ PQ) further boosts results. Overall, we obtain $69$ PQ from GT supervision ($1st$ row) and $62.8$ PQ when using our pseudo-labels ($5th$ row). Remarkably, our labels yield $91\%$ of the GT performance while covering only $14\%$ of the input point cloud (see Tab.˜0.B.1). To further contextualize these results, we note that this evaluation is performed only on classes for which GT instance labels are available. As shown in Fig.˜1, our model learns to segment a much larger variety of classes. Evaluating only the subset of points visible in the camera ($14\%$ of all points, rows 6&76\&7), the gap between the GT-supervised model ($71.8$ PQ) and pseudo-supervised model ($70.7$ PQ) shrinks further. We conclude that by training SAL model using pseudo-labels, we distill the notion of objectness from SAM [31] into our model and obtain segmentation capabilities, similar to the model trained on GT data.
@@
Table 2: Scaling. By contrast to GT data/labels, by increasing the number of queries, SAL improves performance on stuff classes. By increasing the amount of labeled data, we further improve segmentation performance.
@@
$\times$ $24.8$ $32.3$ $66.8$ $29.7$
Semantic distillation.  At inference, SAL predicts a binary segmentation mask and objectness score for each query. To perform zero-shot classification, our model additionally predicts CLIP image tokens. This allows us to not rely on image features during inference but prompt our predicted tokens with arbitrary texts, i.e., object classes. Does the token head loss $\mathcal{L}_{token}$ successfully distill a notion of object semantics from CLIP into our model? To verify this, we perform linear probing experiments: we replace the token distillation head of a trained model with a linear classifier (backbone and instance decoder remain frozen) and train it using GT labels. As can be seen in Tab.˜3, distilling CLIP features into our model significantly improves linear probing performance ($33.1$ PQ, $40$ mIoU) compared to the baseline, not trained with the semantic distillation loss $\mathcal{L}_{token}$ ($20.0$ PQ, $23.4$ mIoU). This confirms that by distilling CLIP features, we inject a notion of semantics into our model, even though our model is not explicitly supervised with any labels containing semantic information – the notion of semantics is learned implicitly via CLIP feature distillation. Without this loss, our zero-shot model ($24.8$ PQ, $29.7$ mIoU) outperforms the linearly probed model tuned using GT semantic labels.
@@ 4.3 Class-Agnostic and Zero-Shot Lidar Panoptic Segmentation
Do we need a vision segmentation foundation model?  To generate a set with a sufficiently high recall, we need to cluster points with varying DBSCAN radius parameters (details in Sec.˜0.A.1). This leads to a large set of predicted masks per scan ($5,413$ avg.). While most of the correct segments are in this set, separating the signal from the noise is difficult. Prior works, such as [52], are limited to moving thing classes as they rely on motion cues to filter static point cloud regions and DBSCAN to group remaining points. Our core insight is that we can utilize vision foundation models that already learned a notion of objectness: SAM generates only $171$ masks per image, reduced to $45$ via flattening, and to $39$ masks after the unprojection, achieving a high recall with only a few instances, or in other words, a high signal-to-noise ratio.
@@
Zero-shot segmentation.  So far, we discussed class-agnostic segmentation. To obtain semantic class predictions per mask in a zero-shot fashion, we devise a simple baseline inspired by [58]. Once masks are unprojected from the image to Lidar space, we utilize associated CLIP features (directly extracted from images using [16]) to perform zero-shot classification via dot product between encoded class prompts and (lifted) CLIP features [67]. This baseline (SAM+DBS+CLIP) can be understood as an approach that directly lifts image features from state-of-the-art image-based models [16] to Lidar. We detail prompts in Sec.˜0.A.3.
@@
There is a significant performance drop between SAM+DBS (replace) baseline, evaluated with Semantic Oracle ($5^{th}$ row, $48.7$ PQ), and zero-shot baseline (SAM+DBS (replace)+CLIP), $7^{th}$ row, $27.5$ PQ. This is not surprising, as zero-shot segmentation is a challenging problem in both image [16] and Lidar domains. To this end, even when transferring features of state-of-the-art image-based models to Lidar we observe a significant performance drop. Remarkably, SAL distills such noisy image-based features to a stronger model ($8^{th}$ row, $33.1$ PQ, +5.6+5.6 PQ). Finally, to emphasize the inherent limitation of baselines, relying on image features, we evaluate on full $360^{\circ}$ point clouds. SAL performs similarly in both settings. The baseline, on the other hand, relies on image features and is, therefore, unable to predict labels out of the camera frustum. This problem can be mitigated by utilizing a setup with denser camera coverage.
@@
Evaluation on super-classes.  We observe that the CLIP classifier often confuses related classes, e.g., car and other vehicle. We additionally evaluate our semantics by prompting with super-classes (as defined in [6]). This alone increases PQ to $63.9$ ($48.5$ when evaluating on full point clouds) and suggests significant potential for improving prompting and text-to-image alignment.
@@
Table 5: Lidar Panoptic Segmentation (LPS) on SemanticKITTI and nuScenes validation sets. We prompt our zero-shot SAL model with the respective class vocabularies and compare its performance to fully-supervised baselines. On SemanticKITTI and nuScenes, we reach $42\%$ and $54\%$ of the fully-supervised model, respectively. This gap reduces significantly when we evaluate super classes.
@@
SAL Zero-shot $24.8$ $32.3$ $66.8$ $17.4$ $30.2$ $28.7$ $48.5$ $59.4$ $78.8$ $52.8$
nuScenes PHNet [36] Full $74.7$ $84.2$ $88.2$ $74.0$ $75.9$ $79.7$
@@
SAL Zero-shot $38.4$ $47.8$ $77.2$ $47.5$ $29.2$ $33.9$ $52.6$ $63.5$ $77.3$ $52.6$